Left Direct Gregorian Technical Observing with IRTC: UT 2009 November 27

Observer: JHill (LBTO)
Telescope Operator: ACeranski
Instrument Support:
Telescope Support: LVaughn
SW support: (Tucson)

Summary

Open for long enough to get a few sets of dice5 offset data through clouds. Dabbled with the script to take field aberration data, but closed by weather before getting a full data set. Open for 4 hours total. Lost 8 hours to weather.

Preparations

Details

00:00 Course #1: Spiral-cut smoked ham, with side of Butch's homemade sausage.

00:20 Open vent doors and start mirror ventilation. There are significant clouds approaching from the SE.

Active Optics have been cleared and AGW homed.

Open

00:52 Open shutter doors to get even more ventilation

00:54 pointto M5_0855 gs=0 TelM="ACQUIRE" PAM- - Adjust IE=+3 CA=+19 IA=-270 at EL=67 deg, AZ=159

01:00 pointto M5_0855 gs=2 TelM="ACTIVE" - guidestar is too faint relative to clouds and twilight.

01:02 pointto M5_0855 gs=4 TelM="ACTIVE" - This star is brighter, so we've got it. Open loop image was 1.4 arcsec FWHM. Active optics improves that to 1 arcsec momentarily, but seeing is quite variable.

01:00 Course #2: Stuffed Cornish Hens, with sides of cranberry sauce and tater tots.

Dice5 cubes on this bright M5 star

01:09 dice5 7 7 DETXY 0.005 500 5 - Seeing is variable between 0.9 and 2 arcsec on guider.

Probe gets stuck in loop 1 position 5
Fri Nov 27 01:16:15.414 2009  Offset alert for left IRTC: AGw command status: timed out waiting for probe to reach destination
Fri Nov 27 01:16:15.466 2009  Offset alert for left IRTC: error: setProbePosition: couldn''t move probe to position (SFP) (x,y,z): 4.24911, 176.419, 18.1756 mm.
Fri Nov 27 01:16:15.517 2009  Offset alert for left IRTC: oacontrol indicated an error moving to new position
Fri Nov 27 01:16:15.568 2009  Offset alert for left IRTC: timed out waiting for probe to reach destination
Fri Nov 27 01:16:15.619 2009  Offset command complete with warnings for left IRTC from IIF

Aaron homes probe, so we are back by loop 2 posiiton 5. That didn't work while offsets still running, so we stop dice5, and do stopAGW/startAGW.

01:28 pointto M5_0855 gs=4 TelM="ACTIVE" - did not receive a valid full CCD readout from guide camera after maximum number of attempts. Giving up.

Stop and restart GCS because it had lost connection to AzCam.

01:31 pointto M5_0855 gs=4 TelM="ACTIVE" -

01:32 dice5 7 7 DETXY 0.005 500 5
ERROR: cannot open `uparm$pipe10987' for writing
  "time | scan (dd, tt, dt) ; dttm  = dd//" "//tt//" "//dt"
     line 39: LBTtools$task_ID.cl (hidden task)

Restart IRAF window.

01:33 dice5 7 7 DETXY 0.005 500 5

The good news: seeing has improved to 0.75 arcsec. The bad news: my IRTC cubes are saturated.

02:09 Notice that the M2 position update is stuck again. Image on guider is 0.67 arcsec and holding focus, so I think the positions are still updating. OSS M2 GUI and PSF M2 GUI both agree on the position that isn't updating, so maybe it is a problem of OSS not updating reflective memory.

02:17 pointto BS9101 TelM="ACQUIRE" gs=0 PAM- No adjustment needed.

02:20 kill and restart OSS.

02:21 pointto BS9101 TelM="ACTIVE" gs=7 - moonlit clouds caused failed acquisition.

02:24 pointto BS9101 TelM="ACTIVE" gs=1 - Star is out of range

etc, etc

02:31 pointto BS9101 TelM="ACTIVE" gs=3 - succeeded with 0.66 arcsec image on guider.

02:37 dice5 7 7 DETXY 1.0 50 5 - through 1-2 mags of moonlit clouds, better by the end
Last single image is irtc.20091127.031124.fits irtc.20091127_00279_cube.fit

03:00 Course 3: Sweet Potato and Apricot soup

03:12 offset 30 30

03:13 takeseq 1 50 (blank sky cube for above dice5) irtc.20091127.031312.fits irtc.20091127_00280_cube.fit

03:15 takeseq 0.005 500 (blank sky cube for previous dice5) irtc.20091127.031505.fits irtc.20091127_00281_cube.fit

Field data on BS9101

Seeing is below 0.6 arcsec, so we'll try some field data. Stars in BS9101 may not be optimally placed.

03:17 cl < home$field_BS9101.cl in /OldRepository/AGW_Data/091127/
Why didn't I see the IDL wfsc images?
Because GCS was still taking WFS images.
(aborted because I forgot to prepdir here)

manually took wfsc 001

03:26 cl < home$field_BS9101.cl
starting with wfsc 002 - 005 on-axis
gs7 wfsc 006 15.41 failed to find the star (Aaron says there is "ick") Why did it only take 1 exposure? Because telescope was not guiding.
on-axis wfsc 007 - 008 failed to find star because of moonlit clouds.

Close

03:45 Close for rising humidity

I created a IDL program "agw_collimate_doit.pro" which is a copy of "agw_collimate.pro" with the DONT_ASK flag set so that it sends Zernikes to PSF without aksing interactively. I updated my Python wrapper "agw_collimate.py" to use that when the -s flag is set.

Open

04:40 Humidity is stable at 87%, and sky is clear. So we reopen. T=1.0 degC, D=-0.9 degC.

04:42 pointto BS9101 TelM="ACTIVE" gs=3 - Found star in acq image, but failed to guide on this very obvious star because of a gradient of moonlight in the field. Background removal should take care of this.

04:47 pointto BS9107 TelM="ACTIVE" gs=70 PA=7.81 - succeeded, but shortly wiped out by clouds.

05:07 cl < home$field_BS9107.cl though clouds
Abort as first acquisition failed.

05:11 pointto BS9107 TelM="ACTIVE" gs=0 - failed because no star in acq image.

05:13 pointto M5_1133 gs=0 TelM="ACQUIRE" PAM- - Adjust IE=-7 CA=24

05:15 pointto BS9107 TelM="ACTIVE" gs=0 - succeeded Seeing is 0.6 arcsec on guider.

Practice Field Data on BS9107

05:20 cl < home$field_BS9107.cl with seeing 0.6 to 0.9 thorugh thin clouds

star wfsc# PA Doug's
on-axis 17-20 0 -466
GS70 21-23 7.81 -475
on-axis 24-25 7.81 -476
GS39 26-28 226.93 -337
on-axis 29-30 226.93 -338
GS97 31-33 124.41 -237
(31 lost to clouds)

on-axis 34-35 124.41 - (acquisition failed for clouds, 34-35 lost)
GS82 36 27.77 (acquisition failed for clouds, 36 lost)>br> on-axis 37-38 27.77 - (failed to see GS in corner, sent bogus Zernikes)
GS35 39-41 315.81
aborted....... we probably measured 3 good stars.

05:52 pointto M5_1133 gs=0 TelM="ACQUIRE" PAM- - Adjust IE=-15 CA=22

Close

05:57 Close for heavy clouds. And that is the end of the night as fog moves in along with the clouds.

-- JohnHill - 19 Nov 2009
I Attachment Action Size Date Who Comment
daq.20091127.gifgif daq.20091127.gif manage 38 K 27 Nov 2009 - 20:48 JohnHill Plot of Telescope and Mirror Temperatures
pmcc.dx.thermal.20091127.gifgif pmcc.dx.thermal.20091127.gif manage 26 K 27 Nov 2009 - 20:49 JohnHill Plot of DX Primary Mirror Temperatures
pmcc.sx.thermal.20091127.gifgif pmcc.sx.thermal.20091127.gif manage 28 K 27 Nov 2009 - 20:48 JohnHill Plot of SX Primary Mirror Temperatures
Topic revision: r5 - 27 Nov 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback